The table below lists the costs and benefits of five mutually exclusive projects.

Project

Costs

Benefits

A

9

12

B

6

5

C

3

4

D

9

11

E

4

6

  1. (6 marks) Calculate the benefit-cost ratio (BCR) of each project:

Project

BCR

A

B

C

D

E

  1. (6 marks) Use incremental benefit-cost analysis to determine the preferred project. Show your work.


The preferred project is Project ___________
Answer #1

a). BCR = Discounted value of benefits/ discounted value of costs

BCRA = 12 / 9 = 1.33

BCRB = 5 / 6 = 0.83

BCRC = 4 / 3 = 1.33

BCRD = 11 / 9 = 1.22

BCRE = 6 / 4 = 1.50

b). The preferred project is Project "E", as its BCR is highest.